# Wetting on Silicone Surfaces

### v1

@@ III.3.2 Elastic
On elastomeric surfaces, the wetting ridge dissipation is more complex [198, 199]. Dissipation in elastomers is non-Newtonian and depends on the (oscillatory) strain rate $\omega$. The crosslinks add (visco)elasticity to the material fabric and the resulting stress is well represented with the Chasset-Thirion model, i.e., $\sigma\propto\epsilon\mu\left(\tau\omega\right)^{m}$ [200]. The moving wetting ridge oscillates with $\omega\propto v/\delta_{\mathrm{e,I}}$. $\tau$ is the rheological relaxation time and $m<1$ yields characteristic shear-thinning of PDMS [201, 202]. These two values vary depending on the degree of crosslinking and the number of free chains in the network [203]. Softer elastomers with more free chains usually have both smaller $\tau$ and smaller $m$. The strain in elastic ridges scales with the ratio of the two elastocapillary lengths $\epsilon\propto\delta_{\mathrm{e,I}}/\delta_{\mathrm{e,II}}$ and the strain rate with $\dot{\epsilon}\propto\epsilon\omega$. Together with the volumetric scaling $V\propto\delta_{\mathrm{e,I}}^{2}$, the dissipative work in the viscoelastic wetting ridge scales as
$P_{\mathrm{diss}}^{\mathrm{vel}}\propto\gamma_{w}vr\alpha^{2}\mathrm{Ca}_{e}^{m},$ (17)
Fig. 6b. Here, we introduce the ”viscoelastic” capillary number $\mathrm{Ca}_{e}=\tau v/\delta_{e,I}$. The thickness of the surface coating affects the shape of the wetting ridge when the length scales of both become comparable [204, 205]. Surface material that accumulates in the elastic wetting ridge cannot be replenished. Thus, the surface dimples around the wetting ridge, in lack of surface material, Fig. 3b. Very thin coating surfaces also limit the size of the wetting ridge as the material is confined (or attached) to an underlying surface. The ridge geometry is in addition to $\delta_{e,\mathrm{I}}$ and $\delta_{e,\mathrm{II}}$ governed by the thin film height. The friction that builds up in the wetting ridge is consequently reduced, as the total dissipation volume, $V_{\mathrm{ridge}}$ decreases [196].
@@ III.3.3 Molecular
In SOCALs, a mobile drop sliding over the surface reorganizes the streched chains in the ridge constantly. While low, the contact angle hysteresis is not zero. This indicates a non-symmetric process: the differing capillary force at the advancing and receding side leads to a non-zero net force as the drop moves. The force that builds up during sliding was measured with a drop friction force instrument [64]. This force scales with the drop speed approximately with $P_{\mathrm{diss}}\propto v^{5/4}$, Fig 6c. While preliminary, these measurements have some important implications: as the variation in synthesis does not appear to influence the effective friction force, therefore, the grafting density and chain length may not play a significant role. However, rigorous investigations are still lacking and decisive conclusions should not be drawn at this stage.
@@ IV Coating Comparison
In this review, we discussed wetting behaviors on PDMS-based surfaces. We discuss how this relatively simple polymer can create a wide design space for surfaces, via polymer-chain configurations, polymerization, and surface anchoring. The complexity in surface manufacturing varies for different surface variants: Creating liquid-infused surfaces without a micro-/nanometric texture is straightforward. It can be achieved by simply spreading PDMS oil on a surface. The process is equally facile for PDMS elastomeric surfaces, but only if no special requirements regarding the coating thickness exist. The ”grafting-from” technique for SOCAL surfaces requires a more elaborate process such as ”chemical-vapor-deposition” (CVD), requiring controlled atmosphere conditions, and sometimes toxic chemicals (e.g. chlorosilanes or alkoxysilanes).
The various surface designs open up a wide variety of concepts and mechanisms. Today, they are increasingly utilized and tuned for a plethora of applications. While similar at first glance, wetting mechanisms such as the wetting ridge develop differently on each PDMS-based surface variant. In the context of a specific application, this brings advantages and disadvantages. For example, self-cleaning are increasingly essential for preserving the energy efficiency of solar panels, bio-contamination reduction of medical equipment, or repelling stains and dirt on outdoor equipment and textiles. For this, all three PDMS-based surface variants are conceptually suitable as they strive to provide drop mobility (low static friction), a feature that is essential for self-cleaning.
Liquid-infused surfaces have freely flowing PDMS oils providing excellent lubrication properties and high drop mobility. An important avenue for future research is to understand how particulate contaminants adheres to the lubricant layer and whether they enter the solid crevices. However, one limitation of liquid-infused surfaces is that lubricant is depleted by the (comparatively) large wetting ridge when they are exposed to harsh environments. Thus, to make these surfaces viable for technological applications, further work is needed to systematically understand the mechanisms of lubricant depletion. Elastomer surfaces, on the other hand, have good material retention and hence mechanical durability, provided by the crosslink network. This advantage, however, brings along higher dissipation margins and declined drop mobility. SOCAL surfaces seem to overcome both challenges, as retention is high due to covalent surface anchors, while dissipation remains low, although not as low as liquid-infused surfaces. This is attributed to a minuscule ridge and the inherent flexibility of molecularly anchored chains that behave in a liquid-like manner. Particulate matter hardly sticks and can be removed easily. However, once the surface experiences critical damage, the nano-thickness of the SOCAL surface can be permanently abraded away. Bulky liquid-infused and elastomer surfaces provide plenty of surface material that - in case of surface damage - can self-heal and increase the surface lifetime.
@@
We currently observe a transition in soft wetting research, shifting from a fundamental understanding perspective to a practical application focus, encompassing fields such as energy and biomedical science. Still, to utilize soft wetting effectively in any application, fundamental understanding is critical, as illustrated in the self-cleaning example, above. This review serves to provide a series of guidelines that link the current state-of-the-art soft wetting knowledge - in particular on silicone surfaces - to mechanistic design features desirable for a myriad of applications.
V Methods and Techniques
V.1 Optical Characterization
V.1.1 Shadowgraphy
@@
The Shadowgraphy method captures shadows of macroscopic objects (e.g., drops of $>1~{}\mathrm{\mu l}$) with high-resolution optical cameras and back-illumination. Shadowgraphs are typically well-contrasted and thus suitable for contour profile extractions of static/dynamic drops using image processing. For example, shadowgraphs are the basis of goniometry, i.e., one of the most common and simplest methods to measure surface wettability. Here, the wetting contact angle is extracted from the drop contour and recovered by various fitting techniques ( e.g. Young-Laplace equation, spherical/elliptical fit, or tangent fit [208, 209]). Goniometry was utilized to demonstrate surface adaptation and dissipation on PDMS gels [44, 210]. The overall shape of the drop contour is governed by the surface tension. If the drop becomes larger than the capillary length, gravity distorts the contour shape from a perfect (hemi)sphere (i.e., oblate sessile drops and prolate pendant drops). When the density of the drop and the prevailing gravity is known, the surface tension can be measured from shadowgraphs (cf. pendant drop measurement). With optically-assisted magnification, microscopic features such as wetting ridges become visually detectable [160]. In place of (vis)light, x-rays were used for high-definition imaging of static wetting ridges [132].
V.1.2 Confocal Microscopy
@@
Confocal microscopy is a non-invasive optical technique that allows us to image samples across both horizontal and vertical planes. Three-dimensional images can also be reconstructed by stacking the two-dimensional horizontal/vertical images. Confocal microscopy has been used to study the static and dynamic properties of drops on liquid-infused surfaces and soft elastomeric surfaces [152, 211, 112, 114]. Fundamentally, confocal microscopy is an extension of fluorescence microscopy where light that is out of focus is filtered out using a pinhole placed at the conjugated point of the optical path. Laser scanning confocal microscopes scan across a sample point by point and combine the scans to produce 3D images with a horizontal resolution of around 200 nm and a vertical resolution of around 1 $\mu$m. The relatively slow point-by-point scan can be accelerated with a spinning archimedean disk. When using confocal microscopy, different fluorescent dyes must be added to the different phases (e.g. drop and coating) being imaged. The use of dyes enables the detection of interfaces between phases of similar refractive index. Typically, inverted confocal microscopes (objective lens below specimen) are used to image drops because this configuration makes it possible to image the contact region between the drop and the surface - which is the region that governs the wetting properties - with minimal optical artifacts. The primary limitation of this technique is that it requires transparent substrates, such as glass.
V.1.3 Interferometry
@@
Interferometric microscopy enables in-plane depth measurements with nano-to-microscopic resolution. In wetting, this has been utilized to directly visualize nanometric features such as contact angles/lines [212, 213], drop profiles [214], wetting ridges [53, 215], and thin films [216, 217, 218, 195, 219]. Coherent light of a “diagnostic beam” travels through the specimen and is brought into superposition with a reference beam with identical coherency. The emerging interference pattern gives the optical path (geometric depth times refractive index) difference between both light beams. The optical path length of the reference beam is calibrated. Different ways of beam calibration brought up various flavors of interference techniques such as Reflection Interference Contrast Microscopy, Digital Holographic Microscopy, or Michelson interferometry, each varying in practical complexity and precision. With information on the refractive index of the specimen, the geometrical depth of the specimen can be determined with a precision of $\leq 10\mathrm{nm}$.
V.1.4 Ellipsometry
@@
Ellipsometry is a non-destructive/non-invasive optical technique that allows the characterization of thin films with sub-Angstrom resolution. The fast scanning speed ($<1$ min for single-wavelength, $<1$ s for multi-wavelength) enables in situ measurements of dynamic films. The technique was used to investigate e.g., precursor spreading of PDMS oil on glass [220, 221, 222] or wetting on SOCAL surfaces [223, 224]. Ellipsometry measures changes in the polarization state of an incident light beam upon interaction with the surface. In ellipsometric measurements, linear polarized light is directed to a film. Depending on the thickness and material properties of the film, the amplitude and phase of the reflected light change due to light-film interactions (absorption/scatter). The light-film interactions are different for the p and s components of the light (parallel and lateral to the plane of the incident), resulting in an altered, elliptical polarization state of the reflected light. The ratio of s/p amplitudes before and after reflection, and the phase shift contains information on the film thickness and optical constants (e.g., refractive index). Ellipsometers either use a single wavelength or multiple wavelengths. In the former case, the angle of incidence is varied and the layer thickness and refractive index are fitted. When working with multiple wavelengths, usually one angle of incidence is used and the film thickness and refractive index are fitted as a function of wavelength. PDMS films on oxidized surfaces (i.e., glass or silicon oxide) are prone to measurement errors as both have almost identical refractive indices.
@@
Figure 8: Different imaging techniques are suitable for different length scales. Most light-based techniques such as (Vis)shadowgraphy or confocal microscopy are bound by the diffraction limit. Interferometry and ellipsometry overcome this limit by exploiting wave superposition and phase shifts, bringing down the resolution limit to sub-nano.
V.2 Mechanical Characterization
V.2.1 Classical Rheology
@@
Rheology describes the stress-strain/shear behavior and the mechanical properties of soft materials such as PDMS. From a wetting perspective, these mechanical properties provide information that can be related to the dynamics and dissipation in the wetting ridge and around the contact line. The term rheology is generic, which includes various geometries and experimental setups to measure dynamic mechanical properties. The most common for silicone-based materials - particularly related to wetting - is parallel-plate shear rheology. In these experiments, silicone-based elastomers, swollen networks, or melts are loaded between two plates. The top plate is typically rotated in a sinusoidal fashion while recording both the stress and the strains. The lower plate can be connected to a Peltier plate, or the entire system can be enclosed into an environmental chamber if temperature or humidity control is needed. Typically, an amplitude strain sweep is conducted at a constant frequency to determine the linear region. A constant strain is then chosen, and a frequency sweep is then conducted to quantify the mechanical properties of the material as a function of frequency (i.e., rate dependence) in the linear regime. Common values that are extracted from the measurement as a function of frequency are the storage $(G^{\prime})$ and loss modulus $(G^{\prime\prime})$. For more liquid-based PDMS (i.e. oils), viscosities can also be measured through a similar method. The storage modulus describes the tendency of the material to store energy (solid-like character) whereas the loss modulus is the tendency to dissipate energy (liquid-like character). Their ratio, $G^{\prime\prime}/G^{\prime}=\tan\delta$ can be associated with damping.
V.2.2 Indenter
@@ V.2.3 Force Microscopy
Forces that build up during drop-sliding on PDMS films can be measured with cantilever set-ups [197]. Such ”force microscopes” were used to measure dissipation on PDMS surfaces of all types. The cantilever is vertically placed over the PDMS surface. Glass capillaries, metal blades, and rods with ring ends have been successfully utilized as cantilevers [197, 235, 236, 196, 237]. The cantilever top is fixed while the lower end hangs freely several micrometers above the surface. Displacements of the lower end - which can be tracked with e.g. shadowgraphy - relate to acting forces and can be extracted with the cantilever spring constant. The spring constant is calibrated by measuring the natural frequency or by applying well-defined loads. To measure dynamic wetting forces on PDMS surfaces, a sessile drop is pushed against the cantilever by moving the surface horizontally. The major drawback of this technique is the invasiveness of the capillary to the drop. It is still not completely understood how the cantilever itself influences the measured friction forces.
@@
Figure 9: Force microscopic measurement conducted on a confocal microscope. Here, $10~{}\mathrm{\mu l}$ water drops slide on on SOCAL surfaces. a) Atto-dyed, steady-state drop footprints, and faintly the probing metal blade. b) Kymograph of the blade deflection upon drop contact, measured with the reflection channel of an $xt$-scan. c) Calibration curve to determine the spring constant of the blade, here approx. $0.21~{}\mathrm{N/m}$.
V.3 Computational Methods
V.3.1 Molecular Dynamics (MD) Simulations

### v2

@@ III.3.2 Elastic
On elastomeric surfaces, the wetting ridge dissipation is more complex [65, 73, 223, 85]. Dissipation in elastomers is non-Newtonian and depends on the (oscillatory) strain rate $\omega$. The crosslinks add (visco)elasticity to the material fabric. For cross-linked PDMS in particular, the resulting material stress can be well described with the power law model proposed by Chasset and Thirion [224], i.e., $\sigma\propto\epsilon G\left(\tau\omega\right)^{m}$. The moving wetting ridge moves with $\omega\propto v/\lambda_{\mathrm{e,I}}$[73]. $\tau$ is the rheological relaxation time and $m<1$ yields characteristic shear-thinning of PDMS [225, 226]. These two values vary depending on the degree of crosslinking and the number of free chains in the network [227]. Softer elastomers with more free chains usually have both smaller $\tau$ and smaller $m$. The strain in elastic ridges scales with the ratio of the two elastocapillary lengths $\epsilon\propto\lambda_{\mathrm{e,I}}/\lambda_{\mathrm{e,II}}$ and the strain rate with $\dot{\epsilon}\propto\epsilon\omega$. Together with the volumetric scaling $V\propto\lambda_{\mathrm{e,I}}^{2}$, the dissipative work in the viscoelastic wetting ridge scales as
$P_{\mathrm{diss}}^{\mathrm{vel}}\propto\gamma_{w}vr\epsilon^{2}\mathrm{Ca}_{e}^{m},$ (17)
Fig. 6b. Here, we introduce the “viscoelastic” capillary number $\mathrm{Ca}_{e}=\tau v/\lambda_{e,I}$. The thickness of the surface coating affects the shape of the wetting ridge when the length scales of both become comparable [69, 75]. Surface material that accumulates in the elastic wetting ridge cannot be replenished. Thus, the surface dimples around the wetting ridge, in lack of surface material, Fig. 3b. Very thin coating surfaces also limit the size of the wetting ridge as the material is confined (or attached) to an underlying surface. The ridge geometry is in addition to $\lambda_{e,\mathrm{I}}$ and $\lambda_{e,\mathrm{II}}$ governed by the thin film height. The friction that builds up in the wetting ridge is consequently reduced, as the total dissipation volume, $V_{\mathrm{ridge}}$ decreases [76].
@@ III.3.3 Molecular
In SOCALs, a mobile drop sliding over the surface reorganizes the stretched chains in the ridge constantly. While low, the contact angle hysteresis is not zero. This indicates a non-symmetric process: the differing capillary force at the advancing and receding side leads to a non-zero net force as the drop moves. The force that builds up during sliding was measured with a drop friction force instrument [222]. This force scales with the drop speed approximately with $P_{\mathrm{diss}}\propto v^{5/4}$, Fig 6c. These measurements have some important implications: as the variation in synthesis does not appear to influence the effective friction force, therefore, the grafting density and chain length may not play a significant role. However, rigorous investigations are still lacking and decisive conclusions should not be drawn at this stage.
@@ IV Coating Comparison
In this review, we discussed wetting behaviors on PDMS-based surfaces. We discuss how this relatively simple polymer can create a wide design space for surfaces, via polymer-chain configurations, polymerization, and surface anchoring. The complexity in surface manufacturing varies for different surface variants: Creating liquid-infused surfaces without a micro-/nanometric texture is straightforward. It can be achieved by simply spreading PDMS oil on a surface. The process is equally facile for PDMS elastomeric surfaces, but only if no special requirements regarding the coating thickness exist. The “grafting-from” technique for SOCAL surfaces requires a more elaborate process such as “chemical-vapor-deposition” (CVD), requiring controlled atmosphere conditions, and sometimes toxic chemicals (e.g. chlorosilanes or alkoxysilanes).
The various surface designs open up a wide variety of concepts and mechanisms. Today, they are increasingly utilized and tuned for a plethora of applications. While similar at first glance, wetting mechanisms such as the wetting ridge develop differently on each PDMS-based surface variant. In the context of a specific application, this brings advantages and disadvantages. For example, self-cleaning is increasingly essential for preserving the energy efficiency of solar panels, bio-contamination reduction of medical equipment, or repelling stains and dirt on outdoor equipment and textiles. For this, all three PDMS-based surface variants are conceptually suitable as they strive to provide drop mobility (low static friction), a feature that is essential for self-cleaning.
Liquid-infused surfaces have freely flowing PDMS oils providing excellent lubrication properties and high drop mobility. An important avenue for future research is to understand how particulate contaminants adhere to the lubricant layer and whether they enter the solid crevices. However, one limitation of liquid-infused surfaces is that lubricant is depleted by the (comparatively) large wetting ridge when they are exposed to harsh environments. Thus, to make these surfaces viable for technological applications, further work is needed to systematically understand the mechanisms of lubricant depletion. Elastomer surfaces, on the other hand, have good material retention and hence mechanical durability, provided by the crosslink network. This advantage, however, brings along higher dissipation margins and declined drop mobility. SOCAL surfaces seem to overcome both challenges, as retention is high due to covalent surface anchors, while dissipation remains low, although not as low as liquid-infused surfaces. This is attributed to a minuscule ridge and the inherent flexibility of molecularly anchored chains that behave in a liquid-like manner. Particulate matter hardly sticks and can be removed easily. However, once the surface experiences critical damage, the nano-thickness of the SOCAL surface can be permanently abraded away. Bulky liquid-infused and elastomer surfaces provide plenty of surface material that - in case of surface damage - can self-heal and increase the surface lifetime.
@@
V Open Questions and Perspectives
PDMS surfaces share several similarities regardless of whether they are in the form of elastomers, SOCALS, or LIS. The wetting ridge is a pivotal element in static and dynamic wetting on each PDMS-based surface type. This is remarkable, considering that the length scale of the wetting ridge can vary by several orders of magnitude and the (macro)molecular configurations can differ significantly between the different variants. It remains an open question whether a unifying model can be developed to describe the wetting dynamics on these different surfaces. To make progress on that front, future studies need to investigate the reorganization of PDMS (oily, crosslinked, or surface-grafted) and the associated dissipation in the dynamic wetting ridge in detail.
The cloaking phenomenon has been observed experimentally on LIS and elastomeric surfaces. Yet, models for the dynamics of cloak formation remain lacking and it remains unknown whether the cloak influences the dissipation mechanisms. Even in equilibrium, the cloak bears an intriguing paradox for water drops: While the spreading coefficient suggests the formation of a stable cloak, dispersive van der Waals (vdW) interactions counterintuitively indicate an attractive disjoining pressure, leading to an unstable film. This paradox may be due to other stabilizing long-range forces (e.g. electrostatic). Further work is needed to understand under which conditions the cloak is stable and whether it influences the dissipation mechanisms. A few considerations and open questions are suggested here for the three types of surfaces.
  • LIS - The PDMS oil in LIS is usually a homogenous, Newtonian, liquid, suggesting a simple design space of the surface. Yet, several methods exist to create the underlying solid substrate and to imbibe the lubricant, leading to a large zoo of LIS designs. Friction laws have been proposed for drops moving on LIS. Still, direct (experimental) validation of the dissipation mechanism in the wetting ridge remains lacking and it is still unclear whether and how the exact geometry of the solid substrate influences the dissipation mechanism. Depletion of lubricant over time remains the biggest challenge that needs to be tackled to design long-lasting PDMS surfaces. The wetting ridge and the cloak are significant causes of depletion when drops slide off lubricated surfaces. As both, the cloak and the wetting ridge originate from thermodynamic forces, avoiding their formation will be challenging. We still lack a quantitative understanding of how much lubricant is depleted per drop which slides off a LIS.
    The prevalence of LIS in the domain of self-cleaning, shear-reduction, and biocompatible surfaces has shown great promise for use in consumable and biotechnological applications. The growth and incubation of cells in LIS surfaces could potentially provide microenvironments for single cell sorting and growth mediums. The unique hierarchical structures of LIS can provide high interfacial interactions for applications requiring cell growth, such as oxygenation or growth media replenishment.
  • Elastomers - The elastomeric ridge is arguably the most complex, owing to the coupling between the liquid and crosslinked PDMS. No other PDMS surface possesses the same degrees of freedom in tunability: crosslink density and distribution of free chains, shear modulus, oil viscosity, oil chemistry, oil-to-network affinity, oil surface tension, and polymer network surface energy. We only recently started to understand the process of oil separation in the wetting ridge, but the consequences of and for sliding drops are still largely unexplored. This is especially evident from a theoretical point of view. What is the interplay between the oil and the crosslinked PDMS in the wetting ridge and how does it change during dynamic wetting? Even for a static scenario, it is still partially unclear what causes separation, or more importantly, what keeps the lubricant trapped inside the network. Answering these question is a key aspect of designing long-term stable elastomeric coatings.
    The formation of cloaks and accumulation of ridges that are composed of PDMS oil remains one of the most severe degradation issues of PDMS-based surfaces. Swelling of PDMS elastomers with liquids that show much less interactions with wetting liquids (e.g. fluorinated oils) compared to PDMS oils (high spreading coefficients) may resolve or slow down the engineering challenge of separation, cloaking, and gradual surface degradation. Further investigation into the use of sustainable infusion liquids or network materials could help facilitate the future development of long-term functional and naturally degradable materials.
  • SOCALS - Wetting ridges on SOCAL are inherently hard to image and investigate, due to the nanometric length scale. While state-of-the-art computation methods are providing a greater understanding of SOCAL surfaces, experimental validation is needed to validate the accuracy of these computational models. This is particularly challenging for water drops. Detailed experimental investigations suffer from the combination of the nanometric length scales and high surface curvatures around the three-phase contact line, caused by contact angles close to 90. To obtain the required resolution, devices with higher sensitivity need to be developed. Alternatively, the problem can be approached using drops with lower liquid polarity, circumventing the challenges accompanied by high curvatures and contact angles. For nonpolar drops, nanometric analytical techniques such as interferometric microscopy (cf. Appendix – Methods and Techniques) might provide sufficient in-depth resolution to resolve the dynamic reorganization of the wetting ridge. A key goal would be to develop constitutive models to describe dissipation on SOCAL surfaces.
    Tuning the silane chemistry on SOCAL surfaces may provide gradients or patterns in glass transition properties, thereby enabling temperature-sensitive surfaces. The use of PDMS may also help shield defects of other organic-material based surfaces owing to the “self-smoothing” nature of highly flexible siloxane chains. Co-integrating different surface chemistries may optimize chain flexibility, giving rise to multi-functional surfaces capable of next-generation easy-to-clean surfaces. Promising applications include on-land or floating photovoltaic applications. More intensive studies are needed to investigate their potential in particulate or dust removal while maintaining high energy throughput during operation.
We are currently observing a transition in the perspective of soft wetting research, shifting from a fundamental understanding to a focus on practical applications, encompassing fields such as energy and biomedical sciences. Overall, PDMS-based surfaces show superior UV-stability, combined with easy-to-clean performance. The integration of PDMS-based surfaces with other organic or inorganic materials may realize useful engineering concepts. For instance, in the emerging field of synthetic biology, PDMS-based microfluidic devices play a crucial role for the synthesis of a broad range of artificial cell precursors. The subsequent assembly of these complex building blocks in functional units remains an ambitious goal of synthetic chemists, physicists and biologists. Fully exploring the potential of PDMS-based surfaces may represent a pathway to improve the synthesis and assembly of cellular building blocks (membranes, proteins, cytoskeleton, etc). Furthermore, the surface energy of PDMS is almost as low as that of environmentally harmful perfluoroalkyls. It may therefore serve as a viable replacement material for many electrode (i.e., Nafion) or membrane materials (i.e. PTFE-based) that are at the core of our energy-water infrastructure.
@@ V Open Questions and Perspectives
To utilize soft wetting effectively in any application, fundamental understanding is still critical, as illustrated in the self-cleaning context (cf. Coating Comparison). This review serves to provide a series of guidelines that link the current state-of-the-art soft wetting knowledge - in particular on silicone surfaces - to mechanistic features desirable for a myriad of applications.
Conflicts of interest
The authors have no conflicts to disclose.
Acknowledgements
Appendix - Methods and Techniques
Optical Characterization
Shadowgraphy
@@ Acknowledgements
The Shadowgraphy method captures shadows of macroscopic objects (e.g., drops of $>1~{}\mathrm{\mu l}$, cf. Fig. 1, 8) with high-resolution optical cameras and back-illumination. Shadowgraphs are typically well-contrasted and thus suitable for contour profile extractions of static/dynamic drops using image processing. For example, shadowgraphs are the basis of goniometry, i.e., one of the most common and simplest methods to measure surface wettability ($\theta$, $\theta_{a}$, $\theta_{r}$, and $\Delta\theta$). Here, the wetting contact angle is extracted from the drop contour and recovered by various fitting techniques ( e.g. Young-Laplace equation, spherical/elliptical fit, or tangent fit [230, 231]). Goniometry was utilized to demonstrate surface adaptation and dissipation on PDMS gels [46, 73]. The overall shape of the drop contour is governed by the surface tension. If the drop becomes larger than the capillary length, gravity distorts the contour shape from a perfect (hemi)sphere (i.e., oblate sessile drops and prolate pendant drops). When the density of the drop and the prevailing gravity is known, the surface tension ($\gamma$) can be measured from shadowgraphs (cf. pendant drop measurement). With optically-assisted magnification, microscopic features such as wetting ridges become visually detectable [186]. In place of (visible) light, x-rays were used for high-definition imaging of static wetting ridges (cf. $\lambda_{e}$) [177], Fig 3b.
Schlieren
Shadowgraphy setups can be upgraded to perform quantitative Schlieren images [232]. For example, wetting-induced deformations of thin PDMS elastomers were measured with Schlieren setups [75]. For this, a collimation lens aligns the illuminating light directly after its source. The parallel light travels through the specimen (i.e., the PDMS surface) and is subsequently focused by a condenser lens. Placing a blade edge at the focal point of the condenser lens blocks the unperturbed light beams. However, any slopes on the surface of the PDMS specimen refract the light and render a visible Schlieren image behind the blade edge. The image intensity of the refracted light and the slope angle of the surface are related and can be found via appropriate calibration.
Confocal Microscopy
@@
Confocal microscopy is a non-invasive optical technique that allows us to image samples across both horizontal and vertical planes. Three-dimensional images can be reconstructed by stacking the two-dimensional horizontal/vertical images. Confocal microscopy has been used to study the static and dynamic properties of drops on liquid-infused surfaces and soft elastomeric surfaces (e.g., $\lambda_{l}$ and $\lambda_{e}$)[68, 233, 81, 82], cf. Fig. 3a, 4ab, 5b inset. Fundamentally, confocal microscopy is an extension of fluorescence microscopy where light that is out of focus is filtered out using a pinhole placed at the conjugated point of the optical path. Laser scanning confocal microscopes scan across a sample point by point and combine the scans to produce 3D images with a horizontal resolution of around 200 nm and a vertical resolution of around 1 $\mu$m, Fig. 8. The relatively slow point-by-point scan can be accelerated with a spinning archimedean disk. When using confocal microscopy, different fluorescent dyes must be added to the different phases (e.g. drop and coating) being imaged. The use of dyes enables the detection of interfaces between phases of similar refractive index. Typically, inverted confocal microscopes (objective lens below specimen) are used to image drops because this configuration makes it possible to image the contact region between the drop and the surface - which is the region that governs the wetting properties - with minimal optical artifacts. The primary limitation of this technique is that it requires transparent substrates, such as glass.
Interferometry
@@
Interferometric microscopy enables in-plane depth measurements with nano-to-microscopic resolution. In wetting, this has been utilized to directly visualize nanometric features such as contact angles/lines ($\theta$) [234, 235], drop profiles [236], wetting ridges ($\lambda_{l}$ and $\lambda_{e}$) [55, 237], and thin films [238, 239, 240, 83, 241]. Coherent light of a “diagnostic beam” travels through the specimen and is brought into superposition with a reference beam with identical coherency. The emerging interference pattern gives the optical path (geometric depth times refractive index) difference between both light beams. The optical path length of the reference beam is calibrated. Different ways of beam calibration brought up various flavors of interference techniques such as Reflection Interference Contrast Microscopy, Digital Holographic Microscopy, or Michelson interferometry, each varying in practical complexity and precision. With information on the refractive index of the specimen, the geometrical depth of the specimen can be determined with a precision of $\leq 10~{}\mathrm{nm}$, Fig. 8.
Ellipsometry
@@
Ellipsometry is a non-destructive/non-invasive optical technique that allows the characterization of thin films with sub-Angstrom resolution. The fast scanning speed ($<1$ min for single-wavelength, $<1$ s for multi-wavelength) enables in situ measurements of dynamic films. The technique was used to investigate e.g., precursor spreading of PDMS oil on glass [242, 243, 244] or wetting on SOCAL surfaces ($H_{B}$) [245, 246]. Ellipsometry measures changes in the polarization state of an incident light beam upon interaction with the surface. In ellipsometric measurements, linear polarized light is directed to a film. Depending on the thickness and material properties of the film, the amplitude and phase of the reflected light change due to light-film interactions (absorption/scatter). The light-film interactions are different for the p and s components of the light (parallel and lateral to the plane of the incident), resulting in an altered, elliptical polarization state of the reflected light. The ratio of s/p amplitudes before and after reflection, and the phase shift contains information on the film thickness and optical constants (e.g., refractive index). Ellipsometers either use a single wavelength or multiple wavelengths. In the former case, the angle of incidence is varied and the layer thickness and refractive index are fitted. When working with multiple wavelengths, usually one angle of incidence is used and the film thickness and refractive index are fitted as a function of wavelength. PDMS films on oxidized surfaces (i.e., glass or silicon oxide) are prone to measurement errors as both have almost identical refractive indices.
@@
Figure 8: Different imaging techniques are suitable for different length scales. Most light-based techniques such as (visible) shadowgraphy or confocal microscopy are bound by the diffraction limit. Interferometry and ellipsometry overcome this limit by exploiting wave superposition and phase shifts, bringing down the resolution limit to sub-nano.
Mechanical Characterization
Classical Rheology
@@
Rheology describes the stress-strain/shear behavior and the mechanical properties of soft materials such as PDMS. From a wetting perspective, these mechanical properties provide information related to the dynamics and dissipation in the wetting ridge and around the contact line. The term rheology is generic, which includes various geometries and experimental setups to measure dynamic mechanical properties. The most common for silicone-based materials - particularly related to wetting - is parallel-plate shear rheology. In these experiments, silicone-based elastomers, swollen networks, or melts are loaded between two plates. The top plate is typically rotated in a sinusoidal fashion while recording both the stress and the strains. The lower plate can be connected to a Peltier plate, or the entire system can be enclosed into an environmental chamber if temperature or humidity control is needed. Typically, an amplitude strain sweep is conducted at a constant frequency to determine the linear region. A constant strain is then chosen, and a frequency sweep is then conducted to quantify the mechanical properties of the material as a function of frequency (i.e., rate dependence) in the linear regime. Common values that are extracted from the measurement as a function of frequency are the storage $(G^{\prime})$ and loss modulus ($G^{\prime\prime}$). For more liquid-based PDMS (i.e. oils), viscosities can also be measured through a similar method. The storage modulus describes the tendency of the material to store energy (solid-like character) whereas the loss modulus is the tendency to dissipate energy (liquid-like character). Their ratio, $G^{\prime\prime}/G^{\prime}=\tan\delta$ can be associated with damping.
Indenter
@@
Forces that build up during drop-sliding on PDMS films can be measured with cantilever set-ups [221]. Such “force microscopes” were used to measure dissipation on PDMS surfaces of all types, Fig. 6. The cantilever is vertically placed over the PDMS surface. Glass capillaries, metal blades, and rods with ring ends have been successfully utilized as cantilevers [221, 257, 258, 76, 259]. The cantilever top is fixed while the lower end hangs freely several micrometers above the surface. Displacements of the lower end (Fig. 9a) - which can be tracked with e.g. shadowgraphy - relate to acting forces and can be extracted with the cantilever spring constant. The spring constant is calibrated by measuring the natural frequency (Fig. 9b) or by applying well-defined loads. To measure dynamic wetting forces on PDMS surfaces, a sessile drop is pushed against the cantilever by moving the surface horizontally. The major drawback of this technique is the invasiveness of the capillary to the drop, 9c. It is still not completely understood how the cantilever itself influences the measured friction forces.
@@
Figure 9: Force microscopic measurement conducted on a confocal microscope. Here, $10~{}\mathrm{\mu l}$ water drops slide on on SOCAL surfaces (cf. Fig. 6c). a) Kymograph of the blade deflection upon drop contact, measured with the reflection channel of an $xt$-scan. b) Calibration curve to determine the spring constant of the blade, here approx. $0.21~{}\mathrm{N/m}$. c) Atto-dyed, steady-state drop footprints, and faintly the probing metal blade. d) overlay of various relaxed drop footprints.
Computational Methods
Molecular Dynamics (MD) Simulations